# Successive Partial Disruptions with Orbital Precession in a White Dwarf–Black Hole System for Repeating GRB 250702B

### v1

@@ Abstract
The peculiar gamma-ray burst GRB 250702B is the longest event ever observed, lasting about one day and exhibiting four prompt-emission flares of $\sim 100$ s with irregular recurrence intervals of at least one hour. To explain this hierarchy of timescales, we consider a scenario in which a stellar object undergoes repeated partial tidal disruptions by a black hole (BH). We find that if a white dwarf (WD) is on a highly eccentric orbit ($e\approx 0.97$) around an intermediate-mass black hole (BH) with $M_{\rm BH}\lesssim 10^{6}\,M_{\odot}$ and $a=50\,R_{\odot}\left(M_{\rm BH}/10^{6}\,M_{\odot}\right)^{1/3}$, the observed properties of GRB 250702B can be naturally reproduced. In this framework, the duration of each flare is determined by the viscous accretion timescale of material stripped near pericenter, with a typical mass $\Delta M\approx 2\times 10^{-2}\,M_{\odot}$. The minimum recurrence time corresponds to the orbital period, while the total activity period is set by the secular orbital evolution timescale leading to the complete disruption of the WD. Furthermore, if $M_{\rm BH}\gtrsim 10^{5}\,M_{\odot}$ and the orbit has a minimum polar angle relative to the BH equatorial plane of $\theta_{\rm min}\gtrsim 0.12~{\rm rad}$, relativistic frame dragging induces $\gtrsim 0.1$ rad precession of the orbital angular momentum between successive pericenter passages, comparable to a typical GRB jet half-opening angle, resulting in intermittent alignment with the observer and irregular flare spacing. The WD experiences $\approx 40$ jet-launch episodes before complete disruption, but only four are expected to be observed on-axis. The remaining off-axis jets become visible at late times, enhancing the radio afterglow by about an order of magnitude, providing a testable prediction of this scenario.
@@ I Introduction
Recently, an exceptionally long-duration gamma-ray burst, GRB 250702B, was detected, exhibiting the longest prompt emission duration ever observed (Neights et al., 2025). The event was first discovered by the Einstein Probe (EP) in X-rays (Li et al., 2025) and was subsequently detected by the Fermi Gamma-ray Burst Monitor (Fermi/GBM) about one day later (Neights et al., 2025). The first gamma-ray flare occurred approximately 10 hours after the initial X-ray detection, followed by several subsequent flares with irregular separations ranging from $\sim$1 to $\sim$25 hours (Li et al., 2025; Zhang et al., 2025; Oganesyan et al., 2025; Neights et al., 2025). Each flare lasts for $\delta t_{\mathrm{flare}}\sim 100$ s. The total activity duration reached $\delta t_{\mathrm{tot}}\sim 1.6$ d, with an isotropic-equivalent luminosity of $L_{\gamma,\mathrm{iso}}\approx 5\times 10^{51}$ erg s-1 at $z=1.036$ (Gompertz et al., 2025). Multiwavelength follow-up observations detected X-ray, infrared, and radio emission (Levan et al., 2025; Carney et al., 2025; O’Connor et al., 2025), and the transient was localized at an offset of 5.7 kpc from the center of its host galaxy (Carney et al., 2025).
The coexistence of short ($\sim 100$ s) flares, irregular hour-scale recurrence, and prolonged day-scale activity makes GRB 250702B highly unusual among known GRBs, suggesting a physical origin distinct from both compact-object mergers and collapsars. Such hierarchical timescales have been proposed to arise in tidal disruption events (TDEs) (Levan et al., 2025; Beniamini et al., 2025; Eyles-Ferris et al., 2025; Li et al., 2025; Granot et al., 2025).
@@ II Physical Picture and Timescale Hierarchy
Figure 1: Our physical picture for GRB 250702B. Panel (a) shows the geodesic orbit of the WD, where the red line represents our numerical result with $a=20\,R_{\odot}$, $e=0.97$, $M_{\rm BH}=1\times 10^{5}\,M_{\odot}$, $a_{\rm spin}=0.9$, and $\theta_{\rm min}=0.12~\mathrm{rad}$ ($7^{\circ}$). The BH spin axis is indicated by the vertical black line. The orbit exhibits relativistic precession. Mass loss from the WD is not included in this calculation. Panel (b) illustrates an unobservable configuration. An accretion disk forms with its angular-momentum vector tilted by an angle $\theta_{\rm d}$ with respect to the BH spin axis. As a result, the jet undergoes Lense–Thirring precession. If $|\theta_{\rm v,BH}-\theta_{\rm d}|>\theta_{\rm j}$, the jet does not intersect the line of sight, and the prompt emission remains unobservable. Panel (c) shows an observable configuration. If $|\theta_{\rm v,BH}-\theta_{\rm d}|\lesssim\theta_{\rm j}$, the jet remains continuously observable throughout the Lense–Thirring precession.
@@ II.1 Repeating Partial Disruptions
We consider a scenario in which a WD of mass $M_{\rm WD}$ and radius $R_{\rm WD}$ is captured by a BH of mass $M_{\rm BH}$ and dimensionless spin parameter $a_{\rm spin}$, most likely as a result of dynamical scattering within its host stellar system. After capture, the WD settles onto an eccentric bound orbit and undergoes a sequence of partial tidal disruption events. During each pericenter passage, a fraction of the WD mass is stripped, remains gravitationally bound to the BH, and is ultimately accreted, while the stellar core survives and continues on a bound orbit. Such repeating partial tidal disruptions have been demonstrated in previous studies (e.g., Zalamea et al., 2010; MacLeod et al., 2014; Chen et al., 2023, 2025). The mass lost during a single pericenter passage can be estimated following Zalamea et al. (2010) as
@@
where $r_{\rm h}\equiv r_{\rm g}\left(1+\sqrt{1-a_{\rm spin}^{2}}\right)$ is the radius of the event horizon of a Kerr BH with $r_{\rm g}=GM_{\rm BH}/c^{2}$. Under these conditions, the stellar core remains gravitationally bound to the BH after each passage and continues on a stable eccentric orbit. As a result, multiple partial tidal disruption events can occur before the WD is fully disrupted. Under this picture, the system naturally produces three hierarchical timescales, $\sim 100$ s, $\sim 1$ h, and $\sim 1$ d, which we associate with the observed properties of GRB 250702B.
@@ II.2.1 Flare duration: viscous timescale
At each pericenter passage, the stripped material is assumed to circularize on a timescale much shorter than the orbital period, forming an accretion disk around the BH. We therefore identify the duration of an individual gamma-ray flare with the viscous timescale of this disk. The circularization radius is taken to be $r_{\rm d}\sim 2r_{\rm p}=2a(1-e)$, where $a$ is the semi-major axis and $e$ is the eccentricity. The viscous timescale is then given by
$\delta t_{\mathrm{flare}}$ $\approx$ $t_{\rm vis}=\frac{1}{h^{2}\alpha}\sqrt{\frac{r_{\rm d}^{3}}{GM_{\rm BH}}}\sim\frac{1}{h^{2}\alpha}\sqrt{\frac{8R_{\rm WD}^{3}}{GM_{\rm WD}\beta^{3}}}$ (4)
$\sim$ $100\left(\frac{\beta}{0.6}\right)^{-\frac{3}{2}}\left(\frac{h}{0.5}\right)^{-2}\left(\frac{\alpha}{0.1}\right)^{-1}$
@@
where $h$ is the disk aspect ratio, $\alpha$ is the viscosity parameter, and $G$ is the gravitational constant. For fiducial accretion-disk parameters $h\sim 0.5$ and $\alpha\sim 0.1$, Eq. (4) implies that achieving the flare duration of $\sim 100~\rm s$ requires the disruption of a compact object, such as the WD.
@@ II.2.2 Minimum flare interval: orbital period
Partial tidal disruptions occur in a nearly periodic manner, with each event taking place at pericenter passage (e.g., Zalamea et al., 2010; MacLeod et al., 2014; Chen et al., 2023, 2025). The characteristic recurrence timescale therefore corresponds to the orbital period, and we identify the minimum flare interval with
$\Delta t_{\rm flare,min}$ $\simeq$ $P_{\rm orb}=2\pi\sqrt{\frac{a^{3}}{GM_{\rm BH}}}$ (5)
$\sim$ $1~\left(\frac{a}{20~R_{\odot}}\right)^{\frac{3}{2}}\left(\frac{M_{\rm BH}}{1\times 10^{5}~M_{\odot}}\right)^{-\frac{1}{2}}~{\rm h}.$
@@ II.2.3 Total activity duration: secular evolution
As the WD loses mass during each pericenter passage, its orbital angular momentum decreases, leading to secular orbital evolution. The characteristic timescale for this evolution can be estimated as
$\delta t_{\mathrm{tot}}$ $\approx$ $t_{\rm sec}\sim\frac{M_{\rm WD}}{\Delta M}P_{\rm orb}$
$\sim$ $1.6\left(\frac{\Delta t_{\rm flare,min}}{1~{\rm h}}\right)\left(\frac{\Delta M}{2\times 10^{-2}~M_{\odot}}\right)^{-1}\left(\frac{M_{\rm WD}}{1~M_{\odot}}\right)~{\rm d}.$
We note that the orbital energy loss per passage is typically much smaller than the angular-momentum loss associated with each partial disruption event (e.g., Zalamea et al., 2010; MacLeod et al., 2014). Consequently, the pericenter distance remains nearly constant over many passages, consistent with recent numerical simulations (e.g., Chen et al., 2023, 2025). Since each passage removes a mass element $\Delta M$, the total number of passages is $\simeq t_{\mathrm{tot}}/\Delta t_{\rm flare,min}\approx 40$. Over these successive passages, the progressive mass loss causes the WD to expand, which increases the tidal radius while the pericenter distance remains fixed. Around the $\sim 40$th passage, the WD undergoes complete tidal disruption, thereby terminating the repeating flare activity.
@@ II.3 Parameter Constraints from Observed Timescales
By combining Eqs. (4), (5), and (II.2.3) with the observed timescale of GRB 250702B, and adopting typical WD parameters of $M_{\rm WD}\sim 1\,M_{\odot}$ and $R_{\rm WD}\sim 0.01\,R_{\odot}$, together with accretion-disk parameters of $h\sim 0.5$ and $\alpha\sim 0.1$, we obtain
$\beta$ $\approx 0.6\left(\cfrac{\delta t_{\rm flare}}{100~{\rm s}}\right)^{-\frac{2}{3}},$ (8)
$\left(\frac{a}{R_{\odot}}\right)^{3}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)^{-1}$ $\approx 0.1\left(\cfrac{\Delta t_{\rm flare,min}}{1~{\rm h}}\right)^{2},$ (9)
$\Delta M$ $\approx 2\times 10^{-2}\,M_{\odot}\left(\cfrac{\Delta t_{\rm flare,min}}{1~{\rm h}}\right)\left(\cfrac{\delta t_{\rm tot}}{1~{\rm d}}\right)^{-1}.$ (10)
The inferred value of $\beta$ satisfies Eq. 2. At this stage, the combination $a^{3}/M_{\rm BH}$ is constrained. Using these constraints, the eccentricity can be estimated from
@@
$\approx$ $0.03\left(\cfrac{\delta t_{\rm flare}}{100~{\rm s}}\right)^{\frac{2}{3}}\left(\cfrac{\Delta t_{\rm flare,min}}{1~{\rm h}}\right)^{-\frac{2}{3}},$
@@
Using the parameters inferred above, we verify whether the system terminates after a finite number of partial disruption events. The mass-loss episodes are expected to occur periodically $\approx 1.6~{\rm d}/1~{\rm h}\approx 40$ times. After $\sim 40$ passages, the WD mass is expected to decrease to of order $\sim 0.1\,M_{\odot}$. Using the mass–radius relation of Zalamea et al. (2010), $R_{\rm WD}=0.013\,R_{\odot}\left(\frac{M_{\rm Ch}}{M_{\rm WD}}\right)^{1/3}\left(1-\frac{M_{\rm WD}}{M_{\rm Ch}}\right)^{0.447}\sim 0.02\,R_{\odot}$, the corresponding tidal radius at the 40th orbit becomes $r_{\rm T}\sim 0.03\,R_{\odot}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)^{1/3}$. Using the orbital parameters inferred from the observed timescales, $(a/R_{\odot})^{3}(M_{\rm BH}/M_{\odot})^{-1}\sim 0.1$, the pericenter distance is $r_{\rm p}\sim 0.02\,R_{\odot}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)^{1/3}$. Since $r_{\rm T}>r_{\rm p}$ at this stage, the WD would undergo complete tidal disruption on the $\sim 40$th passage.

### v2

@@ Abstract
The peculiar gamma-ray burst GRB 250702B is the longest event ever observed, lasting about one day and exhibiting four prompt-emission flares of $\sim 100$ s with irregular recurrence intervals of at least one hour. To explain this hierarchy of timescales, we consider a scenario in which a stellar object undergoes repeated partial tidal disruptions by a black hole (BH). We find that if a white dwarf (WD) is on a highly eccentric orbit ($e\approx 0.97$) around an intermediate-mass BH with $M_{\rm BH}\lesssim 10^{6}\,M_{\odot}$ and $a=60\,R_{\odot}\left(M_{\rm BH}/10^{6}\,M_{\odot}\right)^{1/3}$, the observed properties of GRB 250702B can be reproduced. In this framework, the duration of each flare is determined by the viscous accretion timescale of material stripped near pericenter, with a typical mass $\Delta M\approx 2\times 10^{-2}\,M_{\odot}$. The minimum recurrence time corresponds to the orbital period, while the total activity period is set by the secular orbital evolution timescale leading to the complete disruption of the WD. Furthermore, if $M_{\rm BH}\gtrsim 10^{5}\,M_{\odot}$ and the orbit has a minimum polar angle relative to the BH equatorial plane of $\theta_{\rm min}\gtrsim 0.12~{\rm rad}$, relativistic frame dragging induces $\gtrsim 0.1$ rad precession of the orbital angular momentum between successive pericenter passages, comparable to a typical GRB jet half-opening angle, resulting in intermittent alignment with the observer and irregular flare spacing. The WD experiences $\approx 40$ jet-launch episodes before complete disruption, but only four are expected to be observed on-axis. The remaining off-axis jets become visible at late times, enhancing the radio afterglow by about an order of magnitude, providing a testable prediction of this scenario.
@@ I Introduction
Recently, an exceptionally long-duration gamma-ray burst, GRB 250702B, was detected, exhibiting the longest prompt emission duration ever observed (E. Neights et al., 2025). The event was first discovered by the Einstein Probe (EP) in X-rays (D. Y. Li et al., 2025) and was subsequently detected by the Fermi Gamma-ray Burst Monitor (Fermi/GBM) (E. Neights et al., 2025). The first gamma-ray flare occurred approximately 10 hours after the initial X-ray detection, followed by several subsequent flares with irregular separations ranging from $\sim$1 to $\sim$25 hours (D. Y. Li et al., 2025; J.-P. Zhang et al., 2025; G. Oganesyan et al., 2025; E. Neights et al., 2025). Each flare lasts for $\delta t_{\mathrm{flare,obs}}\sim 100$ s. The total activity duration reached $\delta t_{\mathrm{tot,obs}}\sim 1.6$ d, with an isotropic-equivalent luminosity of $L_{\gamma,\mathrm{iso,obs}}\approx 5\times 10^{51}$ erg s-1 at $z=1.036$ (B. P. Gompertz et al., 2025). Multiwavelength follow-up observations detected X-ray, infrared, and radio emission (A. J. Levan et al., 2025; J. Carney et al., 2025; B. O’Connor et al., 2025), and the transient was localized at an offset of 5.7 kpc from the center of its host galaxy (J. Carney et al., 2025).
The coexistence of short ($\sim 100$ s) flares, irregular hour-scale recurrence, and prolonged day-scale activity makes GRB 250702B highly unusual among known GRBs, suggesting a physical origin distinct from both compact-object mergers and collapsars. Such hierarchical timescales have been proposed to arise in tidal disruption events (TDEs) (A. J. Levan et al., 2025; P. Beniamini et al., 2025; R. A. Eyles-Ferris et al., 2025; D. Y. Li et al., 2025; J. Granot et al., 2025; C. Yuan et al., 2026).
@@ II Physical Picture and Timescale Hierarchy
Figure 1: Our physical picture for GRB 250702B. Panel (a) shows the geodesic orbit of the WD, where the red line represents our numerical result with $a=30\,R_{\odot}$, $e=0.97$, $M_{\rm BH}=1\times 10^{5}\,M_{\odot}$, $a_{\rm spin}=0.9$, and $\theta_{\rm min}=0.12~\mathrm{rad}$ ($7^{\circ}$). The BH spin axis is indicated by the vertical black line. The orbit exhibits relativistic precession. Mass loss from the WD is not included in this calculation. Panel (b) illustrates an unobservable configuration. An accretion disk forms with its angular-momentum vector tilted by an angle $\theta_{\rm d}$ with respect to the BH spin axis. As a result, the jet undergoes Lense–Thirring precession. If $|\theta_{\rm v,BH}-\theta_{\rm d}|>\theta_{\rm j}$, the jet does not intersect the line of sight, and the prompt emission remains unobservable. Panel (c) shows an observable configuration. If $|\theta_{\rm v,BH}-\theta_{\rm d}|\lesssim\theta_{\rm j}$, the jet remains continuously observable throughout the Lense–Thirring precession.
@@ II.1 Repeating Partial Disruptions
We consider a scenario in which a WD of mass $M_{\rm WD}$ and radius $R_{\rm WD}$ is captured by a BH of mass $M_{\rm BH}$ and dimensionless spin parameter $a_{\rm spin}$, most likely as a result of dynamical scattering within its host stellar system. After capture, the WD settles onto an eccentric bound orbit and undergoes a sequence of partial tidal disruption events. During each pericenter passage, a fraction of the WD mass is stripped, remains gravitationally bound to the BH, and is ultimately accreted, while the stellar core survives and continues on a bound orbit. Such repeating partial tidal disruptions have been demonstrated in previous studies (e.g., I. Zalamea et al., 2010; M. MacLeod et al., 2014; J.-H. Chen et al., 2023; J.-P. Chen et al., 2025).
The mass lost during a single pericenter passage can be estimated following I. Zalamea et al. (2010) as
@@
where $r_{\rm h}\equiv r_{\rm g}\left(1+\sqrt{1-a_{\rm spin}^{2}}\right)$ is the radius of the event horizon of a Kerr BH with $r_{\rm g}=GM_{\rm BH}/c^{2}$. Under these conditions, the stellar core remains gravitationally bound to the BH after each passage and continues on a stable eccentric orbit. As a result, multiple partial tidal disruption events can occur before the WD is fully disrupted.
In our picture, as the WD loses mass, its radius increases due to the mass–radius relation (see I. Zalamea et al., 2010), and the expansion of the WD can dominate over the evolution of its Roche lobe, thereby enhancing the degree of Roche lobe overflow and ultimately leading to full disruption. The resulting orbital evolution is assumed to be governed primarily by the secular timescale.
Within this framework, we associate the observed three hierarchical timescales, $\sim 100$ s, $\sim 1$ h, and $\sim 1.6$ d, with the viscous timescale, the orbital period, and the secular evolution timescale, respectively. In the regime $t_{\rm vis}\ll P_{\rm orb}$, angular momentum can in principle be efficiently redistributed through the accretion disk. In the case of nearly conservative mass transfer, this leads to orbital expansion and suppression of further tidal stripping (e.g., A. King, 2022; R. A. Eyles-Ferris et al., 2025). On the other hand, if a fraction of the transferred mass is lost from the system via disk winds and/or jets, the evolution becomes non-conservative, allowing for net angular momentum loss. In our picture, a significant fraction of the transferred mass is accreted onto the BH, while the remaining fraction is expelled via outflows, carrying away a substantial amount of angular momentum. This leads to a non-conservative evolution of the system. In such a regime, the expansion of the WD exceeds that of the Roche-lobe, resulting in progressively enhanced mass stripping [^1] We parameterize $(\dot{J}/J)\sim f(\dot{M}_{\rm WD}/M_{\rm WD})$ and use Eq. (12) of A. King (2022) to determine the condition on $f$ for which the WD expansion rate exceeds that of the Roche-lobe, requiring: $\frac{\dot{R}_{\rm L}}{R_{\rm L}}-\frac{\dot{R}_{\rm WD}}{R_{\rm WD}}=-2\frac{\dot{M}_{\rm WD}}{M_{\rm WD}}\left(\frac{5}{6}+\frac{\zeta}{2}-\frac{M_{\rm WD}}{M_{\rm BH}}\right)+2\frac{\dot{J}}{J}-\frac{\dot{e}}{1+e}<0$. Here, the parameter $\zeta$ is defined by $\dot{R}_{\rm WD}/R_{\rm WD}=\zeta\dot{M}_{\rm WD}/M_{\rm WD}$, and we adopt $\zeta\simeq-1.4$ for a $M_{\rm WD}\sim 1\,M_{\odot}$ WD. Substituting our parameters determined in Section II.3, together with $\dot{M}_{\rm WD}\sim-\frac{\Delta M}{P_{\rm orb}}$ and $\dot{e}\sim-\frac{e}{t_{\rm sec}}$, we obtain the condition $f>0.45$, indicating that angular momentum removal via outflows is required for a non-conservative evolution. In this regime, the relative expansion rate between the Roche-lobe radius and the WD radius evolves on the secular timescale (Eq. (7)).. In the following, we focus on the system in this non-conservative regime.
@@ II.2.1 Flare duration: viscous timescale
At each pericenter passage, the stripped material is assumed to circularize on a timescale much shorter than the orbital period, forming an accretion disk around the BH. We therefore identify the duration of an individual gamma-ray flare with the viscous timescale of this disk. The circularization radius is taken to be $r_{\rm d}\sim 2r_{\rm p}=2a(1-e)$, where $a$ is the semi-major axis and $e$ is the eccentricity. The viscous timescale can be estimated as
$\delta t_{\mathrm{flare,obs}}$ $\approx$ $(1+z)t_{\rm vis}$ (4)
$=$ $\frac{1+z}{h^{2}\alpha}\sqrt{\frac{r_{\rm d}^{3}}{GM_{\rm BH}}}\sim\frac{1+z}{h^{2}\alpha}\sqrt{\frac{8R_{\rm WD}^{3}}{GM_{\rm WD}\beta^{3}}}$
$\sim$ $100~\left(\frac{1+z}{2}\right)\left(\frac{\beta}{0.6}\right)^{-\frac{3}{2}}\left(\frac{h}{0.5}\right)^{-2}\left(\frac{\alpha}{0.3}\right)^{-1}$
@@
where $h$ is the disk aspect ratio, $\alpha$ is the viscosity parameter, and $G$ is the gravitational constant. In the regime considered here, the accretion rate is super-Eddington, which implies an increased aspect ratio (e.g., M. A. Abramowicz et al., 1988; A. Sadowski et al., 2015). We adopt a representative value $h=0.5$. We further use $\alpha=0.3$, consistent with observationally inferred values for hot, fully ionized accretion flows (R. G. Martin et al., 2019). With these parameters, Eq. (4) implies that achieving the flare duration of $\sim 100~\rm s$ requires the disruption of a compact object, such as the WD.
@@ II.2.2 Minimum flare interval: orbital period
Partial tidal disruptions occur at each pericenter passage (e.g., I. Zalamea et al., 2010; M. MacLeod et al., 2014; J.-H. Chen et al., 2023; J.-P. Chen et al., 2025). The minimum flare interval corresponds to the orbital period,
$\Delta t_{\rm flare,min,obs}$ $\simeq$ $(1+z)P_{\rm orb}=2\pi(1+z)\sqrt{\frac{a^{3}}{GM_{\rm BH}}}$
$\sim$ $1~\left(\frac{1+z}{2}\right)\left(\frac{a}{30~R_{\odot}}\right)^{\frac{3}{2}}\left(\frac{M_{\rm BH}}{1\times 10^{5}~M_{\odot}}\right)^{-\frac{1}{2}}~{\rm h}.$
@@ II.2.3 Total activity duration: secular evolution
As the WD loses mass during each pericenter passage, the orbit evolves on a secular timescale. The corresponding secular evolution timescale can be estimated as
$\delta t_{\mathrm{tot,obs}}$ $\approx$ $(1+z)t_{\rm sec}\sim(1+z)\frac{M_{\rm WD}}{\Delta M}P_{\rm orb}$ (7)
$\sim$ $1.6~\left(\frac{1+z}{2}\right)\left(\frac{\Delta t_{\rm flare,min,obs}}{1~{\rm h}}\right)\left(\frac{\Delta M}{2\times 10^{-2}~M_{\odot}}\right)^{-1}$
$\times\left(\frac{M_{\rm WD}}{1~M_{\odot}}\right)~{\rm d}.$
@@ II.3 Parameter Constraints from Observed Timescales
By combining Eqs. (4), (LABEL:eq:Delta_t_flare_min), and (7) with the observed timescale of GRB 250702B, and adopting typical WD parameters of $M_{\rm WD}\sim 1\,M_{\odot}$ and $R_{\rm WD}\sim 0.01\,R_{\odot}$, together with accretion-disk parameters of $h\sim 0.5$ and $\alpha\sim 0.3$, we obtain
$\beta\approx 0.6\left(\cfrac{\delta t_{\rm flare,obs}}{100~{\rm s}}\right)^{-\frac{2}{3}},$ (8)
$\left(\frac{a}{R_{\odot}}\right)^{3}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)^{-1}\approx 0.2\left(\cfrac{\Delta t_{\rm flare,min,obs}}{1~{\rm h}}\right)^{2},$ (9)
$\Delta M\approx 2\times 10^{-2}\,M_{\odot}\left(\cfrac{\Delta t_{\rm flare,min,obs}}{1~{\rm h}}\right)\left(\cfrac{\delta t_{\rm tot,obs}}{1.6~{\rm d}}\right)^{-1}.$ (10)
The inferred value of $\beta$ satisfies Eq. (2). At this stage, the combination $a^{3}/M_{\rm BH}$ is constrained. Using these constraints, the eccentricity can be estimated from
@@
$\approx$ $0.03\left(\cfrac{\delta t_{\rm flare,obs}}{100~{\rm s}}\right)^{\frac{2}{3}}\left(\cfrac{\Delta t_{\rm flare,min,obs}}{1~{\rm h}}\right)^{-\frac{2}{3}},$
@@
Using the parameters inferred above, we verify whether the system terminates after a finite number of partial disruption events. The mass-loss episodes are expected to occur periodically $\approx 1.6~{\rm d}/1~{\rm h}\approx 40$ times. After $\sim 40$ passages, the WD mass is expected to decrease to of order $\sim 0.1\,M_{\odot}$. Using the mass–radius relation of I. Zalamea et al. (2010), $R_{\rm WD}=0.013\,R_{\odot}\left(\frac{M_{\rm Ch}}{M_{\rm WD}}\right)^{1/3}\left(1-\frac{M_{\rm WD}}{M_{\rm Ch}}\right)^{0.447}\sim 0.02\,R_{\odot}$, the corresponding tidal radius at the 40th orbit becomes $r_{\rm T}\sim 0.03\,R_{\odot}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)^{1/3}$. Using the orbital parameters inferred from the observed timescales, $(a/R_{\odot})^{3}(M_{\rm BH}/M_{\odot})^{-1}\sim 0.1$, the pericenter distance is $r_{\rm p}\sim 0.02\,R_{\odot}\left(\frac{M_{\rm BH}}{M_{\odot}}\right)^{1/3}$. Since $r_{\rm T}>r_{\rm p}$ at this stage, the WD would undergo complete tidal disruption on the $\sim 40$th passage.